OpenStudy (anonymous):

maximum value of f(x)=-6x^2+7x+9?

Directrix (directrix):

Are you studying parabolas or derivatives?

OpenStudy (anonymous):

parabolas

OpenStudy (akshay_budhkar):

to find the maximum value find the value of x when\[{d(f(x)) \over dx}=0\], the substitute it in the f(x) to find it

Directrix (directrix):

y = -6x^2+7x+9 The x-coordinate of the vertex is - b /(2a) where a is the coefficient of x^2 and b is the coefficient of the x term. x = -(7) / 2-(6) = 7/12 The corresonding y-value is 6 (7/12)^2 + 7 (7/12) + 9 = 167/24 which is the y-coordinate of the vertex. 167/24 is the maximum value of f(x) See graph at the link: http://www.wolframalpha.com/input/?i=y+%3D+-6x%5E2%2B7x%2B9

Directrix (directrix):

* x = -(7) / 2-(6) = 7/12 should read x = -(7) / 2(-6) = (-7)/ (-12) = 7/12